OpenStudy (anonymous):

Use the Distributive Property to simplify: A.) (x+4)-2(x+5) B.) 1/2(4z+6) C.) (4+5)-(5+2) D.) x(x+7)

OpenStudy (anonymous):

i answered you .. you should use the fact that : a(b+c) = a*b + a*c for example D) x(x+7) = x*x + 7*x = x^2 + 7x